### v1

# Intergenerational Equity of Climate Change Mitigation

Abstract
Today’s decisions on climate change mitigation affect the damage that future generations will bear. Discounting future benefits and costs of climate change mitigation is one of the most important components of assessing efficient climate mitigation pathways. We extend the DICE model with stochastic discount rates to reflect the dynamic nature of discount rates. Stochastic rates give rise to a mitigation strategy, resulting in all model quantities becoming stochastic. We show that the classical calibration of the DICE model induces intergenerational inequality: future generations have to bear higher costs relative to GDP and that this effect worsens under stochastic discount rates. Accounting for additional financing risks, we investigate two modifications of DICE. We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing. We propose a modified optimisation to keep costs below 3% of GDP, resulting in more equal efforts between generations.
Introduction
Climate change is one of the most significant risks of the next centuries 1, as demonstrated by its manifold impacts on society 2. Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide3, a transition to (net) zero emissions is necessary to limit global warming to any constant level. A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world. One traditional method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see 4; 5.
We focus on the DICE model 6; 7 as a simple IAM, which is used to demonstrate the impacts of modifications to the original model due to its simplicity 8; 9; 10; 11; 12. While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points 13; 14; 15; 16; 17; 18; 19, natural feedback processes such as permafrost 9 or abstract catastrophic risks 20; 21. Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimisation function have a strong influence on the resulting optimal pathways 22; 23; 24; 25; 26. Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks 13 or inequalities in the distribution of damages 27.
Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation. Model extensions updating IAMs 11; 12; 19; 28 show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality 29 or energy consumption 30, but also general SCC 31 emphasise the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels. Discount rates, in particular, are an essential component of IAM modelling. Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE 7 versus a normative approach to discounting as in the Stern review 22 continues. The assumptions on discounting have already been discussed in the context of intergenerational 32 and intragenerational inequality 33. Recent studies 34; 35 show that dynamic modelling of discount rates might allow a more precise assessment of the costs of damages and abatement.
We examine the calibration of the DICE model with respect to intergenerational equity. Measuring the costs per GDP of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations. We use an implementation of the DICE model with a flexible time discretisation 36 and couple this model with a comprehensive modelling toolbox for financial markets 37 to enable the integration of valuation methods for complex financial products in IAMs. We add stochastic interest rates to the model and introduce a stochastic abatement policy. The stochastic abatement policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calibration of the model. Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that will significantly improve intergenerational equity: Financing of abatement costs and non-linear discounting of costs, 38. As abatement is a planned investment process, we assume that these costs may be financed by loans with stochastic interest rates. Furthermore, as large projects and investments often incur additional costs 39; 40, we introduce increasing financing costs for large payments via a non-linear discounting. These higher costs for high damages can be viewed as (additional) financing risks. We provide an overview of the extensions in fig. 1. More background is provided in our extensive model description 36.
Basic model DICE-2016-R 7  –  re-implementation with improvements Integrating numerical methods37 (Monte-Carlo simulation, algorithmic differentiation) Improved representation of time-discretisation via an Euler-scheme. Stochastic interest rates g Funding of abatement cost g Non-linear funding of large damage cost Limiting total cost by GDP share
Figure 1: Interest rate related model extensions. We expand our re-implementation of DICE-2016-R with 1) a stochastic interest rate module with an optional stochastic abatement model, 2) the option to allow funding of abatement cost, and, 3) the option to apply non-linear discounting of damage cost. The latter can be used to penalize emission paths that generate cost overruns in terms of cost-per-GDP.
Results
Intergenerational Inequality
The standard DICE model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V/\mathrm{d}C$ and the discount factor.
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (1)
Here $V$ denotes the total welfare, $C_{\mathrm{D}}(t)$ and and $C_{\mathrm{A}}(t)$ denote the damage and abatement cost, respectively, and $T^{\mu=1}$ denotes a parameter representing the abatement policy. The formula presented here is for a reduced model parameterised by the parameter $T^{\mu=1}$, the time by which 100% abatement is achieved.
The criterion eq. 1 is plausible if a single agent optimises its cost but may be problematic if the costs are distributed between different generations. Note that the two individual contributions in eq. 1 occur at different magnitudes at different times, even different generations. fig. 2 shows the temporal distribution of the weighted functions $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$.
Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$. The calibration is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
Hence, marginal gains of one generation are equated to marginal losses of another generation. The calibration does not equalise the burden but the weighted marginal burden. In this sense, the calibration likely introduces intergenerational inequality.
Including stochastic interest rates into DICE
We see in eq. 1 that both the welfare cost sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the numeraire ratio $\frac{N(0)}{N(t)}$ (i.e., the discount factor) are important to determine the intertemporal optimal pathway. Including a stochastic rate model, the numeraire ratio becomes stochastic. Allowing for (stochastic) adaptation of the abatement policy to the interest rate level, all model quantities become stochastic.
We show that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality. fig. 3 depicts the impact of stochastic interest rates, both in average and in a quantile measure, on emissions, optimal abatement, damage and total cost-per-GDP. These quantities are depicted for the classical DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10 % and 90 % quantile values.
Figure 3: Classical DICE (grey) and Stochastic DICE Model (green). Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost). Shaded area shows the 10 % to 90 % percentiles.
Including funding of abatement and funding risks of damage into DICE
We endow the model with two model extensions: First, we modify the abatement costs by allowing these to be funded over time for interest that has to be paid. Second, we include a non-linear dependency of discounting of the damage cost on the magnitude of the cost. This can be viewed as the additional cost associated with financing risk of very large amount 38.
We find that both modifications allow to improve intergenerational equity (in expectation and risk) in the respective optimal emission pathway. fig. 4 shows the effects of the two modifications on the DICE model with stochastic interest rates. Similar results for the DICE model with deterministic interest rates are shown in LABEL:supfig:pathways_deterministic_with_extension.
Figure 4: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
Shaded area shows the 10 % to 90 % percentiles.
Figure 4: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
As shown in fig. 3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future. fig. 3 shows that the trade-off between early abatement costs and later damages can be balanced by the constraint that total costs are limited to a certain share of GDP. This measure is already used to assess national debt 41 and results in a more equal distribution of damages between generations. Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, present generations still face the lowest costs. Appropriate parameters of the non-linear discounting of the damage cost can be used to limit the costs per GDP.
To illustrate the unequal burdens between generations, we use population projections from 2015 to 2100 42 to calculate burdens experienced during the projected lifetime. We measure the burden in percentage-of-GDP, where we choose 3 % as the neutral level (green). While differences are small for the time covered by population projections (fig. 5 and LABEL:supfig:generational_burden_deterministic), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century. Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset. The deterministic model keeps the 3% limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
Figure 5: Illustration of generational inequality: Average cost relative to GDP by year of birth in the different model specifications up to 2200 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
Discussion
We extend the DICE integrated assessment model and compare several extensions for their effect on intergenerational equity. Our analyses apply to integrated assessment models in general and could also be extended to benefit-cost analysis of climate mitigation. We provide a complete open-source implementation of the model illustrating all our extensions, utilising (stochastic) algorithmic differentiation to efficiently calculate sensitivities of model quantities against each other. [^1] The code and data is available at https://gitlab.com/finmath/finmath-climate-nonlinear.
Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP. This is because the calibration does not consider any intergenerational equity but balances the marginal cost changes induced by a policy change. Absolute cost and its distribution over time are irrelevant to the objective function. This effect is driven by the aggregation of discounted welfare into intergenerational welfare. Considering stochastic interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost. This indicates that a repeated re-calibration of the DICE model to current interest rate levels will amplify intergenerational inequality. We use a more applied scenario by considering the funding of abatement cost and non-linear damage costs to show that these additional aspects would increase the incentive for early mitigation, reducing the burden on later generations. More importantly, the non-linear damage cost allow to effectively limit the generational cost to a specific percentage of the GDP.
Using a simple IAM includes several caveats. First, the representation of damages could be improved 43; 11; 12. Second, the lack of regional resolution compared to, e.g., econometric damage estimates 44; 45 ignores regional inequalities in exposure and potential distributional effects of projected costs 46. Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways 31; 47.
Modifying the DICE model in a modular way, we address the issue that the classical model does not consider equity between generations. The proposed extensions accounting for the funding cost of abatement and increased financing cost of large damages can improve this aspect of the intertemporal optimisation. This is amplified when considering stochastic interest rate risk. Constraining total costs of mitigation and damages relative to GDP could be used as a simple heuristic to ensure intergenerational equity of climate mitigation.
Methods
Derivation of intergenerational inequality in the standard model eq. 1
We consider the one-parameter deterministic abatement model
$\mu(t)\ =\ \min\left(\mu(0)+\frac{1-\mu(0)}{T^{\mu=1}}\ t,1.0\right)\text{,}$
where the parameter $T^{\mu=1}$ represents the time for reaching 100% abatement. We note that the calibration of a model with the full parameter set is close to the calibration of the model reduced to one parameter 36. As the calibration is maximising the total discounted value, we find that the optimal parameter $T^{\mu=1}$, representing the equilibrium state of the model, fulfils
$\frac{\mathrm{d}}{\mathrm{d}T^{\mu=1}}\ \int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t$ $\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t$
[…]
where $C_{\mathrm{A}}(t)$ and $C_{\mathrm{D}}(t)$ denote the abatement cost and damage cost, respectively.
Equation 2 allows us to interpret a change in a parameter of the abatement policy as a change in the cash-flow structure of a financial product. Here $\mathrm{d}t(C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t))$ corresponds to the time-$t$ cash-flow, $\frac{\partial V}{\partial C(t)}$ to a weight (like a notional) and $\frac{N(0)}{N(t)}$ to the discount factor.
From eq. 2, we see that the optimisation across generations may result in intergenerational inequality: Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}T^{\mu=1}>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}T^{\mu=1}<0$, the calibration is equating the marginal cost changes of abatement and damage. However, these changes occur at different times, maybe even at different generations. Note that eq. 2 implies eq. 1.
Details on different extensions of DICE-2016-R
We endowed the classical DICE model with several extensions. These are not specific to the DICE model and can also be applied to other IAMs. They are also independent and parameterised in a way that allows to recover the classical model for a specific parameter choice.
Stochastic Discount Rates
The interest rate $r$ is modelled with a classical interest rate model. The results depicted here were produced with a Hull-White model with a constant mean-reversion and volatility. The implementation allows to replace this model with a general term structure model.
Stochastic Abatement Policy
It is well understood that the optimal abatement policy depends on the interest rate level. Hence, postulating stochastic interest rates, the abatement policy should be allowed to adapt to the interest rate level. This can also be interpreted as illustrating the effect of a repeated re-calibration of the model to changed interest rate levels. We assume a simple, functional model that depends on the spot interest rate, namely
$\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot r(t,\omega)\right)\cdot t,1.0\right)\text{.}$ (3)
The simplicity of the models allows us to interpret the parameters. We have that $a_{1}$ in (3) describes the change of the abatement speed by interest rate change,
$a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation.
Funding of Abatement Cost
Let $C_{\mu}(t)$ denote the cost that occur in time $t$ due to the abatement policy $\mu$. In the classical model, the total cost that occur in time $t$ are
$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
where $C_{\mathrm{D}}(t)$ are the damage costs and $C_{\mathrm{A}}(t)=C_{\mu}(t)$ are the abatement costs.
We alter the model by allowing a funding period, such that abatement costs become effective after a period $\Delta T_{\mathrm{A}}$. For the funding, interest has to be paid such that
$C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right).$ (4)
For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, $C_{\mu}=C_{\mathrm{A}}$, and we recover the original model.
Since damages occur instantaneously, we do not consider funding of these. However, we consider a kind of non-linear discounting.
Non-Linear Discounting of Damage Cost
The idea behind non-linear discounting is that the discount factor depends on the level of the amount to be discounted. This can be justified by a model where one has to attribute for the higher default risk of excessively large future cash-flows. The mechanics are that the compensation of this default risk lowers the return (the interest rate) and hence increases the discount factor38.
Put differently, this adds an increasing weight to large or very large damages.
Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e. the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$. We then re-define the effective damage costs as
$C_{\mathrm{D}}(t)\ =\ C^{\circ}_{\mathrm{D}}(t)\cdot DC(C^{\circ}_{\mathrm{D}}(t);t)\text{,}$
where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost of funding large projects.
In our experiment, we can specify this weight as a function of the damage $C_{\circ}(t)$ or as a function of the fraction of the GDP, i.e., $C_{\circ}(t)/GDP(t)$.
The latter approach allows us to penalise damages that exceed a certain percentage of the GDP.
Supplementary Material
Supplementary material is available online.
Code and Data availability
The model code for the stochastic DICE model including the proposed extensions, as well as the experiments generating the data and figures is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
Acknowledgements
We thank Max Singhoff for exploring the calibration of the DICE model with the ADAM solver used in our numerical experiments.
L.Q. received funding from the German Federal Ministry of Education and Research (BMBF) under the research projects QUIDIC (01LP1907A) as well as by Deutsche Gesellschaft für Internationale Zusammenarbeit (GIZ) GmbH on behalf of the Government of the Federal Republic of Germany and Federal Ministry for Economic Cooperation and Development (BMZ).
Author Contribution
C.F. and L.Q. developed the research idea together. C.F. lead the re-implementation of the DICE model and extensions as well as the experiments with input from L.Q. L.Q. lead the plotting implementation with input from C.F. C.F. and L.Q. analysed and interpreted results. C.F. derived the analytic expression for generational inequality. C.F and L.Q. wrote the manuscript.
Competing Interests
The authors declare that they have no competing interests.
Supplementary information for
Accounting for Financing Risks improves
Intergenerational Equity of Climate Change Mitigation
Christian P. Fries and Lennart Quante
[…]
figureSupplementary figures
Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$. This is the nominal sensitivity to abatement policy changes, entering the calibration.
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
supfigure]supfig:pathways_deterministic_with_extension
Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
supfigure]supfig:generational_burden_deterministic
Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$. This is the nominal sensitivity to abatement policy changes, entering the calibration.